Status Report

NASA Space Station On-Orbit Status 20 Sep 2003

By SpaceRef Editor
September 20, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 20 Sep 2003

All ISS systems continue to function nominally, except as noted previously or below.

The crew was off duty most of today, except for some routine tasks, etc.

After wakeup (2:00am EDT), morning chores and breakfast, the crew completed the standard 3-hr. station cleaning that is usually done on Saturday.   [“Uborka stantsii” includes removal of food waste products, cleaning of compartments with vacuum cleaner, wet cleaning of the Service Module (SM) dining table and other surfaces with “Fungistat” disinfectant and cleaning fan screens to avoid temperature rises.]

CDR Yuri Malenchenko terminated the bake-out cycle on the BMP micropurification system’s channel 2, moding the channel back to Purify.  This restored both filter beds to Purify mode.  [The regeneration of the air purifier filter beds is repeated every 20 days.  Each bakeout to space vacuum takes about 24 hours.]

A 30-min. task of FE/SO Ed Lu’s today was to reboot all onboard PCS (portable computer system) machines, a regular weekly maintenance activity.

Ed Lu also prepared the daily IMS “delta” file for database updating, while Malenchenko performed the regular routine maintenance/inspection of the SOZh life support systems in the SM. 

The crew had their the weekly PFCs (private family conferences), via S-band/audio and Ku-band/video..

Ed and Yuri also conducted their the weekly planning conference with the ground via S-band/audio, to discuss next week’s “Look-Ahead Plan” (regularly prepared jointly by MCC-H and MCC-M planners and uplinked ahead of time). 

Both crewmembers worked out with their daily 2.5-h program of physical exercise, on TVIS treadmill, RED expander and VELO cycle with load trainer.

At 6:05am EDT, the crew set up and conducted a 10-min. ham radio session with young amateur radio fans at Kagawa Junior High School in Ube City, Japan.  The school is celebrating its 100-year anniversary this year.  [A list of student questions was uplinked to the crew beforehand.]

Flight attitude is back in LVLH YVV (local vertical/local horizontal — y-axis in velocity vector), i.e. earth-oriented and flying “sidewise”.

The Elektron apparatus was turned on on 9/18 (Thursday) and is producing O2 for the crew (by water electrolysis).  In parallel, air is being supplied from Progress 12P tankage.  [Repressurization is limited (by Flight Rule) not to exceed a max. ppN2 (nitrogen) level of 600 mmHg (which protects for Orlan prebreathe during Russian EVA preparation).]

Weekly Science Update (Expedition Seven — 20th):

This was a very busy week for payloads on ISS, and as usual, the crew did an outstanding job.

GASMAP:   All the data from the last check looks great.  The next health check is not until October.

Human Research Facility/Workstation (HRF WS):   Continuing.

Ultrasound (USND):  Next week the crew will do a scanning session.  Malenchenko will also do a procedure review on Tuesday (9/23) in preparation for the session.

Hand Posture Analyzer (HPA):  The crew had the first HPA session last Wednesday (9/17) and Ed Lu completed all the experiment protocols successfully.  HPA data have been received on  the ground, and data analysis is in progress.  The team is looking forward to watching the videotape of HPA operations.

Pore Formation and Mobility Investigation (PFMI):   The crew successfully completed processing two PFMI samples this week and obtained excellent results. These samples had a small amount of water (0.24%) mixed in with the succinonitrile. Breakdown of the planar interface occurred much earlier after the beginning of growth than occurs on earth. This is because of the greatly reduced convection in the liquid ahead of the interface in microgravity.

Space Acceleration Measurement System (SAMS):  SAMS is operating nominally since being brought back on line.  It is receiving data from four sensors. 

Microgravity Acceleration Measurement System (MAMS):  MAMS continues to measure the quasi-steady acceleration environment using the OSS sensor.  The HiRAP sensor downlink was disabled today after it successfully supported the PFMI operations

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):  Behaving nominally.

Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE):   Current test matrix is complete.

Materials ISS Experiment (MISSE):  In progress.  Deployed outside. Nominal and collecting data.

Cellular Biotechnology Support Systems (CBOSS):   CBOSS-FDI is looking forward to OBT (on-board training) and experiment activities next week.

Earth Knowledge Acquired by Middle School Students (EarthKAM):  Activities are completed for Increment 7.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2):  Activities have been concluded for this Expedition.

Commercial Generic Bioprocessing Apparatus (CGBA):  CGBA is maintaining a nominal temperature of 20°C.  A command was issued to CGBA today to remove old data files in order to maintain proper disk storage space.  GCF (Granada Crystallization Facility) will remain in CGBA until returning on 6S.

Educational Payload Operations (EPO):  It has been a great week for education activities. The video footage has been extremely well received. There is a lot of interest in utilizing it in various NASA education programs and web sites.

Crew Earth Observation (CEO):  A beautifully detailed CEO image of Victoria Falls from the crew’s mapping of the Zambezi River will be published on Earth Observatory this weekend.  Their fine images of Hurricane Isabel were a big hit and in much demand this week, especially from PAO, NOAA, and researchers at the University of Wisconsin.  The researchers hope to correlate ISS/CEO views with data obtained from research aircraft operating concurrently in the storm.  The crew was encouraged to continue to take the excellent images the ground is now starting to receive of the islands, coral reefs, and atolls in the South Pacific.  These are often small, challenging targets to both describe and to spot.  However, near-nadir, long lens views are beautiful to the eye and key to the success of the ongoing international mapping, survey, and inventory of the Earth’s coral reef systems.

Today’s CEO (crew earth observation) targets, taking into account the current LVLH attitude, and including the targets of the Lewis & Clark 200-year memorial locations, were Cape Town, South Africa (looking a touch right for the entire urban region [which stretches well east from the classic city center at the foot of Table Mountain]), Lake Eyre, Australia (nadir pass over the north end of the lake.  Crew was to shoot any water patches within the lake, or in dry lakes beds surrounding the main lake), Amazon smoke (Dynamic event.  Smoke from numerous fires in Rondonia is moving slowly south around the big bend in the central Andes.  Shooting obliques left), Congo-Zimbabwe Biomass Burning (looking for fires and smoke left and right of track.  Also, good sunglint opportunity right of track as the ISS passed up the length of the long Lake Malawi), Patagonian Glaciers (looking right of track for detailed views.  Glacier tongues at low altitude are sensitive to temperature change and are widely quoted as indicators of multiyear climate change), Recife, Brazil (nadir pass over this early colonial port city [one of the closest to Portugal]), Ed’s Hill, N. I., New Zealand (looking right of track on North Island’s east coast), and Chao lava, Chile (this unusual lava mass [black and highly wrinkled due to very high viscosity] is at nadir.  Shooting a series of detailed views under track should have captured this feature).
CEO images can be viewed at the websites

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 4:00am EDT).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On, 16A (air also being supplied by Progress 12P).  Vozdukh CO2 scrubber is On (in Manual Mode 5/3).  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is operating.  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On; SKV-2 is Off.

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Directed Position mode (blind/triple-angle mode for drag reduction, non solar-tracking).
  • SM batteries:  Battery #7 is disconnected; battery #4 is in “Cycle” mode; all other batteries (6) are in “Partial Charge” mode (battery #1 is degraded). 
  • FGB batteries:  Battery #6 is disconnected; battery #4 is in “Cycle” mode; all other batteries (4) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-2 MDM is Off; PL-1 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Propulsion System (PS):

  • Total propellant load available: 3922 kg (8646 lb) as of 9/18  [SM(755) + FGB(2815) + Progress M(352) + Progress M-1(0)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH -YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, -y-axis in velocity vector [yaw: 90.2 deg, pitch: -10.8 deg, roll: -3 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • GB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-2 is prime).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at MBS PDGF #1 with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 7:52am EDT [= epoch]):